OpenStudy (openstudy9):

WILL MEDAL! (x-8)(x-k)=x^2-5xk+m In the equation above, k and m are constants. If the equation is true for all values of x, what is the of m? A. 8 B.16 C.24 D.32 E.40

OpenStudy (tkhunny):

What was wrong with the last post with no work shown? It is no different from this post with no work shown. 1) Don't post with no work shown. 2) Have you considered expanding the LHS and looking at paired coefficients on the RHS?

OpenStudy (xmissalycatx):

Try plugging in your values paired with your answer choices to see what works. :P

OpenStudy (anonymous):

Let's solve for m. (x−8)(x−k)=x2−5xk+m Step 1: Flip the equation. −5kx+x2+m=−kx+x2+8k−8x Step 2: Add -x^2 to both sides. −5kx+x2+m+−x2=−kx+x2+8k−8x+−x2 −5kx+m=−kx+8k−8x Step 3: Add 5kx to both sides. −5kx+m+5kx=−kx+8k−8x+5kx m=4kx+8k−8x Answer: m=4kx+8k−8x

OpenStudy (anonymous):

you can probably figure out the rest :D Hope that helped